# Exploring the Mechanisms Behind Non Aromatic Fluorescence with Density Functional Tight Binding Method

### 20 Jan 2024 v2

@@ 4.4.2 L-Glutamine
Figure 9: Oscillator strength distribution for the transition $S_{1}\to S_{0}$ for L-gln and L-pyro-amm systems, employing the unit cell in the QM/MM simulations. The inset shows the same distribution at higher oscillator strength.
@@ 5 Conclusions
In this study, we focused on exploring the phenomenon of Non-Aromatic Fluorescence (NAF) in L-pyro-amm and its comparison with the non fluorescent L-gln system, leveraging the computational efficiency offered by Density Functional Tight-Binding Theory (DFTB). Our application of DFTB theory yielded a robust structural and chemical description of the hydrogen bonds (HB) present in these structures. The correct behavior of the HBs in the L-gln system, as well as the characteristic double-well structure in Short Hydrogen Bonds (SHB) present in L-pyro-amm, was successfully reproduced. These properties were recognized for their pivotal role in NAF across various systems.
@@
The integration of Trajectory Surface Hopping (TSH) with TD-DFTB in our study provided a mutual consensus of results with higher-level theories, exemplifying the versatility and efficacy of this combined approach. The computational efficiency inherent in DFTB enabled us to attain additional statistical significance, particularly advantageous in comprehending the less-frequent non-radiative decay in the L-pyro-amm fluorescence system. Despite these strengths, subtle differences surfaced. Agreement in modes like carbonyl and proton transfer with our previous work highlighted the consistency of key features. On the other hand, modes such as the de-planarization display more differences compared to the previous TD-DFT findings. The precise origin of this observed difference could be an inherent limitation of the DFTB method or arise as a consequence of our non adiabatic dynamics approach, specifically the utilization of the LZ method. In this method, the focus in determining non radiative probabilities is on the energy gap between the states. As soon as the system reaches a minimum energy gap, the algorithm initiates a transition between potential energy surfaces. In contrast, FSSH method incorporates non adiabatic coupling vectors, which significantly influence the hopping criteria, taking into account both the energy gap and the vectors. Consequently, our results exhibit a slightly accelerated decay to the ground state with LZ compared to those in FSSH, reducing the likelihood of observing the minor contribution from the deplanarization mode. Clearly, this discrepancy necessitates further investigation in future studies.
@@
In addition, our investigation extended to larger systems within the QM regions and incorporating environmental effects through the QM/MM approach, reveals the presence of a consistent non-radiative mechanism in the L-gln system. The scalability of our simulations to encompass more degrees of freedom in a larger system elucidates that the observed decay mechanisms persist but manifest over extended time. The computational efficiency afforded by DFTB method proves indispensable in capturing these prolonged processes, which would be impractical with traditional ab-initio methods. In the case of L-pyro-amm, the introduction of environmental factors induces a discernible blue-shift in the absorption/excitation spectra. Despite this, a detailed analysis between ground and excited states simulations allows for the faithful reproduction of the experimentally observed stoke-shift in this system.
@@
GDM, DB and AH thank to the European Commission for funding on the ERC Grant HyBOP 101043272. GDM, UNM and AH also acknowledges to CINECA supercomputing for the resource allocation (project NAFAA-HP10B4ZBB2 and V-CoIns-HP10BY0AET).

### 09 Apr 2024 v3

@@ 4.4.2 L-Glutamine
Figure 9: Oscillator strength distribution for the transition $S_{1}\to S_{0}$ for L-gln (blue dashed line) and L-pyro-amm (red solid line) systems, employing the unit cell in the QM/MM simulations. The inset shows the same distribution at higher oscillator strength.
@@ 5 Conclusions
In this study, we explored the phenomenon of Non-Aromatic Fluorescence (NAF) in L-pyro-amm and its comparison with the non fluorescent L-gln system, leveraging the computational efficiency offered by the density functional tight-binding (DFTB) method. Our application of DFTB yielded a robust structural and chemical description of the hydrogen bonds (HB) present in these structures. The correct behavior of the HBs in the L-gln system, as well as the characteristic double-well structure in short hydrogen bonds present in L-pyro-amm, was successfully reproduced. These properties were recognized as being pivotal for NAF across various systems.
@@
The integration of trajectory surface hopping (TSH) with time-dependent DFTB in our study provided a mutual consensus of results with higher-level theories, exemplifying the versatility and efficacy of this combined approach. The computational efficiency inherent in TD-DFTB enabled us to attain additional statistical significance, particularly advantageous in comprehending the less-frequent non-radiative decay in the L-pyro-amm fluorescence system. Despite these strengths, subtle differences surfaced. Agreement in modes like carbonyl and proton transfer with our previous work highlighted the consistency of key features. On the other hand, modes such as the de-planarization display more differences compared to the previous time-dependent density functional theory findings. The precise origin of this observed difference could be an inherent limitation of the DFTB method or arise as a consequence of our non adiabatic dynamics approach, specifically the utilization of the Landau-Zener method (LZSH). In this method, the focus in determining non radiative probabilities is on the energy gap between the states. As soon as the system reaches a minimum energy gap, the algorithm initiates a transition between potential energy surfaces. In contrast, Fewest-Switches method (FSSH) incorporates non adiabatic coupling vectors, which significantly influence the hopping criteria, taking into account both the energy gap and the vectors. Consequently, our results exhibit a slightly accelerated decay to the ground state with LZSH compared to those in FSSH, reducing the likelihood of observing the minor contribution from the deplanarization mode. Clearly, this discrepancy necessitates further investigation in future studies.
@@
In addition, our investigation extended to larger systems within the QM regions and incorporating environmental effects through the QM/MM approach, reveals that the absolute values for the QM/MM absorption spectra do not precisely match experimental observations. This discrepancy can potentially be mitigated by employing MM polarizable models72 or increasing the size of the QM region73, 74. Despite these differences, our results present a consistent non-radiative mechanism in the L-gln system. Futhermore, the scalability of our simulations to encompass more degrees of freedom in a larger system elucidates that the observed decay mechanisms persist but manifest over extended time. The computational efficiency afforded by DFTB method proves indispensable in capturing these prolonged processes, which would be impractical with traditional ab-initio methods.
6 Associated Content
[…]
@@ 6 Associated Content
The authors would like to express their gratitude to Bálint Aradi for engaging in insightful discussions and giving feedback regarding the research. GDM, DB and AH thank to the European Commission for funding on the ERC Grant HyBOP 101043272. GDM, UNM and AH also acknowledges to CINECA supercomputing for the resource allocation (project NAFAA-HP10B4ZBB2 and V-CoIns-HP10BY0AET). CRL-M and MAS acknowledge financial support from the German Research Foundation (DFG) through Grant No. FR 2833/82-1. CRL-M extends special thanks to UNM for the invitation to the ”Workshop on Frontiers in Excited State Electronic Structure Methods: from Spectroscopy to Photochemistry” last year in Trieste, which served as the catalyst for initiating this collaboration between our groups.